OpenStudy (kpoprocks21):

Find the number of possible negative real zeros for f(x)=x^3-4x^2-3x-9. Please help:)

OpenStudy (solomonzelman):

\(\color{#000000}{ \displaystyle f(x)=3x^3+4x^2-5x +24 }\) would have possible zeros of, \(\color{#000000}{ \displaystyle \pm 24/3 }\) \(\color{#000000}{ \displaystyle \pm 8/3 }\) \(\color{#000000}{ \displaystyle \pm 6/3 }\) \(\color{#000000}{ \displaystyle \pm 4/3 }\) \(\color{#000000}{ \displaystyle \pm 3/3 }\) \(\color{#000000}{ \displaystyle \pm 2/3 }\) \(\color{#000000}{ \displaystyle \pm 1/3 }\)

OpenStudy (solomonzelman):

basically the constant term (and all of it's divisors) divided by the first coefficient.